# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 3.1 Conjugacy properties of sue distributions in multivariate linear models
which yields, by Corollary 1, a $p$–variate Student’s $t$ posterior for ${\bm{\beta}}$. This result provides an important finding which clarifies that, in specific contexts of potential practical interest, Student’s $t$ – Student’s $t$ conjugacy can be attained, thereby expanding some earlier findings in Song and Xia [57] on a simpler formulation. As is clear from the expression of the likelihood, for this property to hold it is necessary to incorporate the classical location dependence on ${\bm{\beta}}$ via ${\bf{X}}{\bm{\beta}}$, together with a weak form of additional dependence induced by the scaling term $\alpha_{\bm{\beta}}=[\nu+({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})]/(\nu+p)$. Recalling Zhang et al. [65], under weakly informative Student’s $t$ priors employed in practice for ${\bm{\beta}}$, such an effect tends to be small, and most of the dependence between ${\bm{\beta}}$ and ${\bf{y}}$ is through the classical linear predictor ${\bf{X}}{\bm{\beta}}$. In addition, notice that, when ${\bm{\beta}}\sim\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu)$, then $({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})/p$ has $F$ distribution with degrees of freedom $p$ and $\nu$, which implies that for moderate $p$ and $\nu$, the term $({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})/p$ and, hence $\alpha_{\bm{\beta}}$, shrink around $1$.
Section 3.2, shows that the conjugacy properties of sue distributions derived in Proposition 1 extend even beyond multivariate linear models for continuous response vectors, to cover, in particular, also generalizations of multivariate probit and multinomial probit under elliptical or skew–elliptical link functions.
@@ 3.2 Conjugacy properties of sue distributions in multivariate binary models
${\bf{y}}=[1(\bar{y}_{1}>0),\ldots,1(\bar{y}_{n}>0)]^{\top},\quad\mbox{with}\quad\bar{{\bf{y}}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}},$ (23)
where $1(\cdot)$ is the indicator function, $\bar{{\bf{y}}}=(\bar{y}_{1},\ldots,\bar{y}_{n})^{\top}\in\mathbb{R}^{n}$ and $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$. Proposition 2 clarifies from a general perspective that sue conjugacy can be established also in these contexts. This result unifies and extends contributions by Durante [32], Fasano and Durante [36], Anceschi et al. [4] and Zhang et al. [65] on particular sue sub–classes; i.e., suns and specific skew–elliptical distributions in the sue family; see also Remark 3.
@@
Consider the binary random vector ${\bf{y}}\in\{0,1\}^{n}$, defined as in (23) and let ${\bm{D}}_{\bf{y}}=\mbox{diag}(2y_{1}-1,\ldots,2y_{n}-1)$. Moreover, assume again $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$ with parameters partitioned as in (21). Then, $({\bm{\beta}}^{\top},\bar{{\bf{y}}}^{\top})^{\top}$ is a sue with dimensions $(p+n,q)$ and parameters defined as in Equation (22). In addition
  1. 1.
@@
    ${\bm{\Pi}}_{{\bf{y}}\mid{\bm{\beta}}}=\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle F_{n+q}\left(\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},g^{(n+q)}_{Q_{\bm{\beta}}({\bm{\beta}})}\right)}{\displaystyle F_{q}(\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(q)}_{Q_{\bm{\beta}}({\bm{\beta}})})},\qquad\mbox{for all}\ {\bf{y}}\in\{0,1\}^{n},$
    and parameters available in closed form according to the following equations
@@
$\begin{bmatrix}{\bm{\beta}}\\ \bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\end{bmatrix}={\bf{A}}_{\bf{y}}\begin{bmatrix}{\bm{\beta}}\\ \bar{{\bf{y}}}\end{bmatrix},\quad\mbox{with}\quad{\bf{A}}_{\bf{y}}=\begin{bmatrix}{\bf{I}}_{p}&\bm{0}\\ \bm{0}&{\bm{D}}_{\bf{y}}\end{bmatrix},$ (24)
where $({\bm{\beta}}^{\top},\bar{{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with parameters as in Equation (22). Therefore, by Lemma 1, we have that $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bf{A}}_{\bf{y}}{\bm{\xi}}^{\dagger},{\bf{A}}_{\bf{y}}\bm{\Omega}^{\dagger}{\bf{A}}^{\top}_{\bf{y}},{\bm{\Delta}}_{{\bf{A}}_{\bf{y}}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with
@@
Under the above construction, the prior for ${\bm{\beta}}$ follows directly by the closure under marginalization of sue distributions.
@@
As for the likelihood of ${\bf{y}}$, recall that $\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\mathbb{P}(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}>\bm{0}\mid{\bm{\beta}})$. In addition, notice that by applying the results in Proposition 1 to the random vector $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}$, we have $(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}({\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(n+q)}_{Q_{\bm{\beta}}(\bm{{\bm{\beta}}})}),$ with parameters defined as in Proposition 2. Therefore, $\mathbb{P}({\bf{y}}\mid{\bm{\beta}})$ coincides with the cumulative distribution function, evaluated at $\bm{0}$, of the sue random variable $(-\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}(-{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},-{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(n+q)}_{Q_{\bm{\beta}}(\bm{{\bm{\beta}}})})$. Hence, by applying Equation (4) to such a sue yields
@@
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle F_{n+q}\left(\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},g^{(n+q)}_{Q_{\bm{\beta}}({\bm{\beta}})}\right)}{\displaystyle F_{q}(\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(q)}_{Q_{\bm{\beta}}({\bm{\beta}})})},\quad\mbox{for all}\ {\bf{y}}\in\{0,1\}^{n},$
@@
as in Proposition 2. To conclude the proof of Proposition 2, note that $({\bm{\beta}}\mid{\bf{y}})$ is distributed as $({\bm{\beta}}\mid\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}>\bm{0})$. As a result, the posterior distribution follows directly by applying Lemma 3 to the sue random vector $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}$. ∎
@@
Proposition 2 clarifies that sue distributions possess fundamental conjugacy properties also when combined with specific models for multivariate binary data. Such a result extends the one recently derived by Zhang et al. [65] under model (23) with a specific focus on a skew–elliptical joint distribution for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ which enforces lack of correlation between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ while inducing an elliptical prior for ${\bm{\beta}}$. This construction can be derived, under simple linear algebra operations, as a particular case of the general sue assumption for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ in Proposition 2, which crucially allows to recover more general Bayesian formulations, including priors beyond the symmetric elliptical family. Such a connection with the contribution by Zhang et al. [65] is helpful to showcase the practical impact of extending conjugacy to broader classes of models beyond classical multivariate and multinomial probit. Examples 34 further stress this aspect with a focus on sun and sut distributions.
Example 3 (Multivariate unified skew–normal (sun) conjugacy).
@@
A direct and natural strategy to adapt the model studied in Example 1 within the binary data context, is to consider $y_{i}=1(\bar{y}_{i}>0)$ with $(\bar{y}_{i}\mid{\bm{\beta}})\sim{\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$, independently for $i\in\{1,\ldots,n\}$, and ${\bm{\beta}}\sim\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$. Such a model is studied in the Supplementary Materials of Anceschi et al. [4] as a broad extension of classical probit models to skewed link functions, further facilitating generalizations to multivariate and multinomial binary responses. Leveraging standard properties of multivariate Gaussian cumulative distribution functions, the resulting likelihood in Anceschi et al. [4] can be alternatively re–expressed as proportional to $\Phi_{2n}([({\bm{D}}_{y}{\bf{X}}{\bm{\beta}})^{\top},\bm{0}^{\top}]^{\top};\bm{\Sigma})$, where $\bm{\Sigma}$ is a block matrix partitioned as $\bm{\Sigma}_{11}=\sigma^{2}{\bf{I}}_{n}$, $\bm{\Sigma}_{21}=\bm{\Sigma}^{\top}_{12}={\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}$ and $\bm{\Sigma}_{22}={\bf{I}}_{n}$.
@@
To recast the above Bayesian formulation within those studied in Proposition 2, consider again the setting $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$ with parameters partitioned as in Example 1. This assumption, combined with model (23) and the proof of Proposition 2, implies $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q+n}({\bf{A}}_{\bf{y}}{\bm{\xi}}^{\dagger},{\bf{A}}_{\bf{y}}\bm{\Omega}^{\dagger}{\bf{A}}^{\top}_{\bf{y}},{\bm{\Delta}}_{{\bf{A}}_{\bf{y}}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+n+q)}),$ with

### v3

@@ 3.1 Conjugacy properties of SUE distributions in multivariate linear models
which yields, by Corollary 1, a $p$-variate Student’s $t$ posterior for ${\bm{\beta}}$. This result provides an important finding which clarifies that, in specific contexts of potential practical interest, Student’s $t$ – Student’s $t$ conjugacy can be attained, thereby expanding some earlier findings in Song and Xia [57] on a simpler formulation. As is clear from the expression of the likelihood, for this property to hold it is necessary to incorporate the classical location dependence on ${\bm{\beta}}$ via ${\bf{X}}{\bm{\beta}}$, together with a weak form of additional dependence induced by the scaling term $\alpha_{\bm{\beta}}=[\nu+({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})]/(\nu+p)$. Recalling Zhang et al. [65], under weakly informative Student’s $t$ priors employed in practice for ${\bm{\beta}}$, such an effect tends to be small, and most of the dependence between ${\bm{\beta}}$ and ${\bf{y}}$ is through the classical linear predictor ${\bf{X}}{\bm{\beta}}$. In addition, notice that, when ${\bm{\beta}}\sim\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu)$, then $({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})/p$ has $F$ distribution with degrees of freedom $p$ and $\nu$, which implies that for moderate $p$ and $\nu$, the term $({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})/p$ and, hence $\alpha_{\bm{\beta}}$, shrink around $1$.
Section 3.2, shows that the conjugacy properties of SUE distributions derived in Proposition 1 extend even beyond multivariate linear models for continuous response vectors, to cover, in particular, also generalizations of multivariate probit and multinomial probit under elliptical or skew-elliptical link functions.
@@ 3.2 Conjugacy properties of SUE distributions in multivariate binary models
${\bf{y}}=[1(\bar{y}_{1}>0),\ldots,1(\bar{y}_{n}>0)]^{\top},\qquad\qquad\bar{{\bf{y}}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}},$ (23)
where $1(\cdot)$ is the indicator function, $\bar{{\bf{y}}}=(\bar{y}_{1},\ldots,\bar{y}_{n})^{\top}\in\mathbb{R}^{n}$ and $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$. Proposition 2 clarifies from a general perspective that SUE conjugacy can be established also in these contexts. This unifies and extends contributions by Durante [32], Fasano and Durante [36], Anceschi et al. [4] and Zhang et al. [65] on particular SUE sub-classes; i.e., SUNs and specific skew-elliptical distributions in the SUE family; see also Remark 3.
@@
Consider the binary random vector ${\bf{y}}\in\{0,1\}^{n}$, defined as in (23) and let ${\bm{D}}_{\bf{y}}=\mbox{diag}(2y_{1}-1,\ldots,2y_{n}-1)$. Moreover, assume again $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$ with parameters partitioned as in (21). Then, $({\bm{\beta}}^{\top},\bar{{\bf{y}}}^{\top})^{\top}$ is a SUE with dimensions $(p+n,q)$ and parameters defined as in (22). In addition
  1. (a)
@@
    ${\bm{\Pi}}_{{\bf{y}}\mid{\bm{\beta}}}=\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle F_{n+q}\left(\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},g^{(n+q)}_{Q_{\bm{\beta}}({\bm{\beta}})}\right)}{\displaystyle F_{q}(\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(q)}_{Q_{\bm{\beta}}({\bm{\beta}})})},$
    for all ${\bf{y}}\in\{0,1\}^{n}$, and parameters available in closed form according to the following equations
@@
$\begin{bmatrix}{\bm{\beta}}\\ \bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\end{bmatrix}={\bf{A}}_{\bf{y}}\begin{bmatrix}{\bm{\beta}}\\ \bar{{\bf{y}}}\end{bmatrix},\qquad\qquad{\bf{A}}_{\bf{y}}=\begin{bmatrix}{\bf{I}}_{p}&\bm{0}\\ \bm{0}&{\bm{D}}_{\bf{y}}\end{bmatrix},$ (24)
where $({\bm{\beta}}^{\top},\bar{{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with parameters as in (22). Therefore, by Lemma 1, we have that $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bf{A}}_{\bf{y}}{\bm{\xi}}^{\dagger},{\bf{A}}_{\bf{y}}\bm{\Omega}^{\dagger}{\bf{A}}^{\top}_{\bf{y}},{\bm{\Delta}}_{{\bf{A}}_{\bf{y}}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with
@@
Under the above construction, the prior for ${\bm{\beta}}$ follows directly by the SUE closure under marginalization.
@@
As for the likelihood of ${\bf{y}}$, recall that $\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\mathbb{P}(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}>\bm{0}\mid{\bm{\beta}})$. Moreover, by applying the results in Proposition 1 to the random vector $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}$, we have $(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}({\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(n+q)}_{Q_{\bm{\beta}}(\bm{{\bm{\beta}}})}),$ with parameters defined as in Proposition 2. Therefore, $\mathbb{P}({\bf{y}}\mid{\bm{\beta}})$ coincides with the cumulative distribution function, evaluated at $\bm{0}$, of the SUE random variable $(-\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}(-{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},-{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(n+q)}_{Q_{\bm{\beta}}(\bm{{\bm{\beta}}})})$. As a consequence, by applying (4) to such a SUE yields
@@
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle F_{n+q}\left(\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},g^{(n+q)}_{Q_{\bm{\beta}}({\bm{\beta}})}\right)}{\displaystyle F_{q}(\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},g^{(q)}_{Q_{\bm{\beta}}({\bm{\beta}})})},$
@@
for all ${\bf{y}}\in\{0,1\}^{n}$, as in Proposition 2. To conclude the proof, note that $({\bm{\beta}}\mid{\bf{y}})$ is distributed as $({\bm{\beta}}\mid\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}>\bm{0})$. As a result, the posterior distribution follows directly by applying Lemma 3 to the SUE random vector $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}$. ∎
@@
Proposition 2 clarifies that SUE distributions possess fundamental conjugacy properties also when combined with specific models for multivariate binary data. This result extends the one recently derived by Zhang et al. [65] under model (23) with a specific focus on a skew-elliptical joint distribution for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ which enforces lack of correlation between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ while inducing an elliptical prior for ${\bm{\beta}}$. Such a construction can be derived, under simple linear algebra operations, as a particular case of the general SUE assumption for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ in Proposition 2, which crucially allows to recover more general Bayesian formulations, including priors beyond the symmetric elliptical family. This connection with the contribution by Zhang et al. [65] is helpful to showcase the practical impact of extending conjugacy to broader classes of models beyond classical multivariate and multinomial probit. Examples 34 further stress this aspect with a focus on SUN and SUT distributions.
Example 3 (Multivariate unified skew-normal (SUN) conjugacy).
@@
A direct and natural strategy to adapt the model studied in Example 1 within the binary data context, is to consider $y_{i}=1(\bar{y}_{i}>0)$ with $(\bar{y}_{i}\mid{\bm{\beta}})\sim{\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$, independently for $i\in\{1,\ldots,n\}$, and ${\bm{\beta}}\sim\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$. Such a model is studied in the supplementary materials of Anceschi et al. [4] as a broad extension of classical probit models to skewed link functions, further facilitating generalizations to multivariate and multinomial binary responses. Leveraging standard properties of multivariate Gaussian cumulative distribution functions, the resulting likelihood in Anceschi et al. [4] can be alternatively re-expressed as proportional to $\Phi_{2n}([({\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}})^{\top},\bm{0}^{\top}]^{\top};\bm{\Sigma})$, where $\bm{\Sigma}$ is a block matrix partitioned as $\bm{\Sigma}_{11}=\sigma^{2}{\bf{I}}_{n}$, $\bm{\Sigma}_{21}=\bm{\Sigma}^{\top}_{12}={\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}$ and $\bm{\Sigma}_{22}={\bf{I}}_{n}$.
@@
To recast the above Bayesian formulation within those studied in Proposition 2, consider the setting $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$ with parameters partitioned as in Example 1. This assumption, combined with model (23) and the proof of Proposition 2, implies $({\bm{\beta}}^{\top},\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n,q+n}({\bf{A}}_{\bf{y}}{\bm{\xi}}^{\dagger},{\bf{A}}_{\bf{y}}\bm{\Omega}^{\dagger}{\bf{A}}^{\top}_{\bf{y}},{\bm{\Delta}}_{{\bf{A}}_{\bf{y}}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}}),$ with